### v2

# LLMs Are In-Context Bandit Reinforcement Learners

Abstract
Large Language Models (LLMs) excel at in-context learning (ICL), a supervised learning technique that relies on adding annotated examples to the model context. We investigate a contextual bandit version of in-context reinforcement learning (ICRL), where models learn in-context, online, from external reward, instead of supervised data. We show that LLMs effectively demonstrate such learning, and provide a detailed study of the phenomena, experimenting with challenging classification tasks and models of sizes from 500M to 70B parameters. This includes identifying and addressing the instability of the process, demonstrating learning with both semantic and abstract labels, and showing scaling trends. Our findings highlight ICRL capabilities in LLMs, while also underscoring fundamental limitations in their implicit reasoning about errors.
Machine Learning, ICML
[…]
kbpurple
1 Introduction
[…]
Figure 1: Illustration of in-context bandit reinforcement learning. The context shows a sequence of user queries, model responses, and feedback in the Banking77 77-label classification domain. The model learns in-context from rewards given to its previous predictions. The final prediction (shown in red) represents the model’s current guess.
Large language models (LLMs) have been shown to exhibit in-context learning (ICL), a form of supervised learning that does not require parameter updates (Brown et al., 2020). ICL relies on including supervised input-output pairs in the LLM context (i.e., prompt), [^1] We use the terms prompt and context interchangeably. and it has proven effective with both few (Brown et al., 2020) and many (Bertsch et al., 2024; Agarwal et al., 2024) examples. We ask whether the ability to learn in-context extends to contextual bandit reinforcement learning (RL), i.e., whether language models can effectively perform in-context reinforcement learning (ICRL) with stateful single-step interaction episodes. This is a desirable property of LLMs, as they can learn new tasks interactively, in an online setting, at deployment time, without requiring demonstration data, but based on rewards on their outputs. This learning signal may be human-generated, automatically provided (i.e., a program successfully completes; Gehring et al., 2024), or even AI-generated (Zhang et al., 2024a).
ICRL naturally combines ICL and reinforcement learning (RL). In contrast to ICL, as an RL process, ICRL does not rely on annotated labels or a fixed dataset. Instead of constructing the LLM context from supervised input-output pairs, the LLM context is constructed from triplets of an input, a model’s predicted output, and its reward. As more input examples are observed, the model has access to additional triplets, leading to an online and continual learning scenario, where model capabilities improve over time. These triplets are followed by a new input, for which the model must predict an output. Figure 1 illustrates bandit ICRL prompting.
We study the ICRL capabilities of Llama 3.1 (Llama Team, 2024), Qwen2.5 (Qwen et al., 2024), and Gemini 1.5 Flash (Gemini Team, 2024) [^2] We also conduct early experiments on Phi-3.5-mini (Abdin et al., 2024), included in the appendix., using several many-label classification benchmarks, with a focus on contextual bandit RL (i.e., stateful single-step episodes).
We find that LLMs demonstrate innate ICRL capabilities, but that two choices are critical for effective learning. First, a relatively high stochasticity level is needed to encourage exploration. Second, using only triples with positive rewards performs best, corresponding to how the classic REINFORCE algorithm (Williams, 1992) essentially ignores zero-reward examples. The latter choice creates a cosmetic similarity to ICL. However, ICRL remains fundamentally different: in ICRL the model must actively explore to find (i.e., generate) these positive triplet pairs, rather than having an expert annotator provide them.
A recurring observation in our experiments is the relative instability of the process, as performance can suddenly dip significantly, before often quickly recovering. We propose a new method, Stochastic ICRL, to add stochasticity to the prompt construction by only sampling some of the past episodes observed in context, instead of increasing the temperature. This enhances exploration, stabilizes performance, and maintains relatively shorter contexts. Interestingly, this also allows the model to learn in the presence of negative signals as well, but still less effectively than without.
Finally, we observe the scaling trends of ICRL, by studying all Qwen2.5 modeling sizes between 0.5B and 72B. As expected, small models exhibit limited capabilities, while larger variants deliver the strongest performance.
Overall, we show that LLMs display ICRL abilities that lead to consistent and often substantial gains across all tasks we study, in both the naive and stochastic versions. For example, in the Banking77 (Casanueva et al., 2020) classification task, Qwen2.5-7B improves from 6.2% zero-shot accuracy to 72.2% through Stochastic ICRL. Our code, data, and experimental logs are available at https://lil-lab.github.io/icrl/.
2 In-context Reinforcement Learning
ICL (Brown et al., 2020) operates by providing a model with annotated demonstrations of a task. A demonstration includes an input (e.g., What is the best football club in Europe?) and its corresponding annotated output (e.g., AC Milan). ICL’s reliance on pre-existing gold-standard labeled data follows the common supervised learning paradigm, although without any change in the model parameters.
ICRL follows the reinforcement learning paradigm (Sutton & Barto, 2018), where models learn by reinforcing their own good behaviors and suppressing their own bad choices. Instead of providing models with correct demonstrations, the model generates an output given an input, then observe the outcome (i.e., reward) of its prediction. It learns from the reward signals, in an online learning setting, all within the context. In this study we focus on a contextual bandit RL scenario, a restricted variant of RL, where the length of each episode is one step.
Formally, the model $\pi$ observes an input $x^{(t)}\sim\mathcal{D}$ sampled from the data distribution $\mathcal{D}$ at time $t$, generates a prediction $\hat{y}^{(t)}$, and then observes a reward $r^{(t)}\sim R(x^{(t)},\hat{y}^{(t)})$. We denote the tuple $(x^{(t)},\hat{y}^{(t)},r^{(t)})$ as an episode. This formulation does not assume access to datasets of correct demonstrations, but to a reward (i.e., feedback) function.
In common RL terminology, the model $\pi$ is the policy, the input $x^{(t)}$ is the state, [^3] In the bandit literature, the state is often called context, and hence the name contextual bandits. We do not use this term to avoid confusion with the LLM context. and the prediction $y^{(t)}$ is the action. Throughout our formulation, the policy is also conditioned on previous episodes in the form of an LLM context, similar to how supervised examples are provided in ICL. These past episodes are not part of the RL state. Instead, the context is used to perform in-context policy improvement, similar to how past episodes are used to perform policy improvement in conventional RL (e.g., via parameter updates).
We design several methods to elicit ICRL from LLMs. The Naive approach is a straightforward implementation of ICRL following the common ICL recipe (Section 2.1). The Stochastic approach (Section 2.2) is an alternative to increasing sampling temperature, but with more stability. In Appendix B.3, we propose Approximate ICRL, an additional approach that shares similarities with Stochastic, while being more efficient in high-memory setups.
2.1 Naive and Naive+ ICRL
Algorithm 1 Naive and Naive+ ICRL
0:
0:  $\mathcal{D}$: Data distribution
0:  $\pi$: Language model policy
0:  $R$: Reward function
0:
1:  Init buffer $\mathcal{E}\leftarrow\emptyset$
2:  for $t=1,2,3,\dots$ do
[…]
4:    Observe input $x^{(t)}\sim\mathcal{D}$
5:    Sample prediction $\hat{y}^{(t)}\sim\pi(\cdot\mid C,x^{(t)})$
6:    Observe reward $r^{(t)}\sim R(x^{(t)},\hat{y}^{(t)})$
7:    if $r^{(t)}\leq 0$  then
8:      Continue to next $t$ Only in Naive+ ICRL
9:    end if
10:    Add episode to buffer $\mathcal{E}\leftarrow\mathcal{E}\cup\{(x^{(t)},\hat{y^{(t)}},r^{(t)})\}$
11:  end for
Algorithm 1 describes the Naive approach, as well as Naive+, a variant that only considers examples with positive reward. Omitting lines 78 gives Naive ICRL, the most straightforward way to implement ICRL. At each time step $t$, the model observes a new example $x^{(t)}$, produces a prediction $\hat{y}^{(t)}$, and receives a reward $r^{(t)}$. Every such model interaction creates an episode, which is appended to the buffer $\mathcal{E}$. For each new interaction, we construct a context $C$ from prior episodes (line 3). In Naive, this context is simply all past episodes. Naive+ ICRL adds lines 78 and ignores negative-reward episodes, retaining only positive episodes in the buffer. In both variants, as the LLM’s context window becomes full, we drop older episodes to maintain a sliding window of the most recent interactions.
Empirically, Naive does not learn effectively (Section 4; Figure 2), while Naive+ is very effective, especially with relatively high sampling temperature. The gap between the two indicates the failure of Naive is due to the presence of examples with negative reward.
Critically, even when only using positive examples, ICRL still differs from supervised ICL in that it relies on the model’s generations rather than a fixed set of annotated demonstrations. This much more challenging scenario necessitates the model to explore and iteratively refine its outputs through reinforcement; without this capability, further learning does not occur.
2.2 Stochastic ICRL
Stochastic ICRL utilizes model sensitivity to prompt composition as an avenue to increase exploration, instead of the increased temperature of Naive. Changes in prompt composition have been widely observed to lead to variance in LLM behavior, including through changes in the set of ICL examples (Zhang et al., 2022; Liu et al., 2022; Chen et al., 2023; Levy et al., 2023), seemingly meaningless stylistic changes (Sclar et al., 2024; Lu et al., 2022), or even interventions based on entropy in the embedding space (Rahn et al., 2024). Generally, this property of LLMs is not viewed positively. However, it introduces stochasticity into the ICRL process, and thereby introduces exploration.
Stochastic introduces context stochasticity by randomly choosing the subset of past episodes to include in the prompt each time the model observes a new input. As in Naive+ ICRL, Stochastic only uses episodes with positive reward, as we empirically observe it to lead to better results.
Algorithm 2 Stochastic ICRL
0:
0:  $\mathcal{D}$: Data distribution
0:  $\pi$: Language model policy
0:  $R$: Reward function
0:  $p_{\text{keep}}$: Prob. to keep examples in context
0:
1:  Init episode buffer $\mathcal{E}\leftarrow\emptyset$
2:  for $t=1,2,3,\dots$ do
3:    Init empty context $C^{(t)}\leftarrow[\,]$
4:    for $e\in\mathcal{E}$ do
5:      $b\sim\text{Bernoulli}(p_{\text{keep}})$
6:      if $b=1$ then
7:         Add episode to context $C^{(t)}\mathrel{+}=e$
8:      end if
9:    end for
10:    Observe input $x^{(t)}\sim\mathcal{D}$
11:    Sample prediction $\hat{y}^{(t)}\sim\pi(\cdot|C^{(t)},x^{(t)})$
12:    Observe reward $r^{(t)}\sim R(x^{(t)},\hat{y}^{(t)})$
13:    if $r^{(t)}>0$ then
14:      Add episode to buffer     $\mathcal{E}\leftarrow\mathcal{E}\cup\{(x^{(t)},\hat{y^{(t)}},r^{(t)})\}$
15:    end if
16:  end for
Algorithm 2 describes Stochastic ICRL. For each input, we construct a new context (lines 39). We decide what past episodes to include in this context by sampling from a Bernoulli variable parameterized by $p_{\text{keep}}$ (lines 49). We sample independently for each past episode. This results in different implicit reasoning for each input, because each is done with a different context. When storing past episodes, we only include episodes with positive reward (lines 1315).
With small $p_{\text{keep}}$, Stochastic will encounter the issue of the LLM context window saturating much later than Naive. However, deploy ICRL for enough interactions, and the context window will saturate, even for models with the largest windows. Similar to naive, we downsample the context if it overflows the LLM context window. We do it by removing episodes from the sampled $C^{(t)}$ uniformly at random until the context fits the model’s context window.
3 Experimental Setup
Models
We use the instruction-tuned versions of Llama 3.1 8B  (Llama Team, 2024) and Qwen2.5 (Qwen et al., 2024) for all model sizes. [^4] We include in the appendix early experiments with Phi-3.5-mini (Abdin et al., 2024), which generally performs worse due to relative model weakness. For the hardest tasks, we also experiment with Gemini 1.5 Flash 8B (Gemini Team, 2024). [^5] We limits our experiments with Gemini due to costs. Overall, we spent \$2,120 USD on Gemini API calls. We use all models in a multi-turn chat format. We compute the maximum number of episodes the context window can take for each model-task combination (Appendix C.2). We use constrained decoding to generate model predictions, as in recent work on ICL (Bertsch et al., 2024).
Tasks
We follow Bertsch et al.’s (2024) study of many-shot ICL in focusing on five classification problems: Banking77 (77 labels; Casanueva et al. (2020)), CLINC150 (150 labels; Larson et al. (2019), NLU (68 labels; Liu et al. (2021)), TREC (6 labels; Li & Roth (2002); Hovy et al. (2001)), and TREC-fine (50 labels; Li & Roth (2002); Hovy et al. (2001)). Because of the large output spaces (up to 150 labels in CLINC150), these tasks are particularly challenging for large language models, as empirically shown by Bertsch et al. (2024) and replicated in our zero-short results. The classification problem creates a contextual bandit scenario. The labels in each dataset are used to compute rewards, and are never shown to the model. Appendix C.3 provides additional information about the datasets.
The datasets are of different sizes. The size of the datasets dictates the number of time steps in our experiments. We randomly sub-sample Banking77, CLINC150, and NLU to 10k examples. TREC and TREC-fine are smaller, so we only use 5k training examples for each. This allows the experiments to be of relatively standard length. The training data corresponds to the data distribution $\mathcal{D}$ in our algorithms. We also sub-sample all test sets to 500 examples each, to reduce the computational cost of experiments. NLU does not provide a standard test set, so we create our own train and test splits. In all experiments, the datasets contain the same examples in the same order.
Semantic vs. Abstract Class Names
We study using both the original class names and abstract labels. The original class names carry important semantic information, which gives the model helpful clues on how input examples map to them (e.g., the output class name calendar update in CLINC150 is a strong hint to which input queries may apply to it). Experiments with abstract labels remove this information by mapping all labels to meaningless abstract strings (e.g., label5). Experiments and results use the original (semantic) labels by default, unless noted explicitly that they are using abstract labels.
Rewards and Prompt Design
We simulate interactive binary rewards from perfect human actors interacting with the system. We do so by comparing the model outputs with the gold label of each input. This is a common practice in studies of the effect of rewards on learning processes (Gao et al., 2023; Lightman et al., 2024), for practical convenience. We deterministically transform the numerical rewards into a natural language format indicating if the model prediction is correct or not, which is more suitable for LLM reasoning. Appendix C.1 provides more details about our prompting.
Evaluation
We report running test accuracy, using the held-out test set of each dataset. We compute it every 500 steps for each test example separately, using the context used to process that step’s training example. In the appendix, we also report regret, the forgone utility from an actual model prediction in comparison to the oracle choice.
Comparisons
We compare ICRL with the zero-shot setting, which corresponds to the performance on the test set without any in-context examples. [^6] We visually highlight the zero-shot performance with sampling temperature $T=1.0$, which is the temperature we use for Naive and Stochastic experiments. Zero-shot accuracy with Naive+’s temperature of $2.0$ can still be observed by looking at the accuracy at the first step of the Naive+ curves. We also report a supervised ICL upper bound by testing performance with the maximum possible number of gold-standard supervised demonstrations in context. These results use expert demonstrations, which the ICRL results have no access to in our learning process. As expected these ICL results outperform the ICRL trends we report. However, the reliance on annotated demonstrations makes them not comparable to the ICRL results. We provide them to get an idea of the upper bound of ICL in these scenarios.
Figure 2: Performance of ICRL. Naive, Naive+, and Stochastic held-out test results for Llama and Qwen and all tasks. Naive+ and Stochastic consistently outperform zero-shot (i.e., first step) and Naive, while also showing consistent trends of continual improvement as more data is observed. 12(b) in Appendix D details start and end accuracies.
4 Results and Analysis
Figure 2 shows the test accuracies for Llama 3.1 8B and Qwen2.5 7B. As an upper bound to in-context learning, we also show the performance of an oracle with access to the gold labels for the maximum number of in-context examples that the model can fit. Unless specified otherwise, we use $p_{keep}=0.1$ and sampling temperature $T=1.0$ for Stochastic and zero-shot, $T=1.0$ for Naive, and $T=2.0$ for Naive+. We choose the best parameters for each ICRL method and include our analysis in Appendix B.
Figure 3: Reward ablations. Test accuracies of Naive and Stochastic with different reward signals. Positive reward only is the best choice for both methods. With Naive, no other strategy facilitates learning. 3(b) in Appendix D details start and end accuracies.
LLMs Learn In-Context From Their Own Predictions and Rewards
Both Naive+ and Stochastic effectively learn in all tasks and for both models, showing significant improvements over zero-shot. Naive+ and Stochastic improve over the zero-shot accuracies by between 28.6–74.4% for Llama, and 29.2–68.4% for Qwen. In general, accuracies approach the upper bound in many settings, demonstrating the strong bandit ICRL capabilities of Llama and Qwen. For both models, both approaches also demonstrates a monotonic growth in performance over time (Stochastic especially), suggesting that with more data performance would improve further. This is especially evident for the most challenging datasets, i.e., the ones with the most labels (i.e., Banking77, CLINC150, NLU), as they generally require more exploration to learn how to map the inputs to the large number of potential outputs.
[…]
Reward Signals Are Crucial, but Mistakes Remain Challenging
Figure 3 shows ablations studies. Removing rewards or inverting them brings about negligible gains over zero-shot performance for both Naive and Stochastic models. This demonstrates that learning is driven by the reward signals, and not simply by the inclusion of domain examples in the context (i.e., domain effect; Min et al., 2022; Pan et al., 2023; Lyu et al., 2023; Kossen et al., 2024).
Unlike Naive, which only performs effectively when exclusively positive-reward episodes are considered (i.e., Naive+), Stochastic partially maintains its learning capabilities even when exposed to negative outcomes. Although its performance is negatively impacted, this suggests that our stochastic approach prevents the model from becoming overwhelmed by signals it struggles to interpret. Notably, Stochastic remains robust even when 10% of the rewards are inverted (i.e., noisy), indicating resilience to noise, which is likely in human-feedback settings.
Overall, our ablations highlight two key points: (1) LLMs can learn online from their predictions only when a reward signal is present, and (2) LLMs exhibit inherent limitations in implicitly learning from mistakes (i.e., without explicit reasoning, as in Wei et al.’s (2022) Chain-of-Thought).
(a) ICRL Scaling Trends. We evaluate Qwen models from 500M to 70B parameters using both Naive+ and Stochastic. Performance improves substantially over zero-shot accuracy for all sizes, but smaller models plateau at lower accuracies, indicating that ICRL performance correlates positively with model size. 5(g) in Appendix D details start and end accuracies.
(b) Stability of Naive+ and Stochastic ICRL. We measure stability by computing Spearman’s rank correlation between accuracy and time step. Except for TREC and TREC-fine (which give inconclusive results), Stochastic exhibits more stable learning on Banking77, CLINC150, and NLU. Larger models also show higher stability, mirroring trends in (a). 6(g) in Appendix D details $\rho$ values.
Figure 5: Comparison of Qwen models (500M–72B). We analyze scaling accuracy gains (a) and stability differences (b).
Label Semantics Contribute, but ICRL Occurs Without It
[…]
Figure 4 summarizes our findings on Llama 8B, Qwen2.5 7B, and Gemini 1.5 Flash 8B. To contextualize the results, we include upper bound results from standard supervised ICL with as many gold demonstrations (with abstract labels) as the context can handle, which generally succeeds across tasks (except for a lower performance on TREC-fine). With just one exemplar per label, ICRL nearly reaches the upper bound for all tasks and models. We also stress that just including one gold demonstration per label is not always enough to reach good performance, and the online process is still important: for example, Llama with exemplars, when tested on CLINC150 with Stochastic, reaches non-negligeble accuracies only after 3k steps. In the absence of any exemplars, the ICRL process still manages to build informative contexts, though overall accuracy is not surprisingly lower. For instance, Qwen and Llama achieve higher than 45% accuracy on NLU and TREC with both Naive+ and Stochastic, indicating that the models learn and refine their output mappings over time. In contrast, Gemini excels when the correct mapping is provided but struggles significantly in a purely exploratory setting.
Given the domain effect observed with semantic labels, and the relatively low, even if significant results observed with abstract labels, the question arises if this learning is due to a domain effect. For example, Llama on CLINC150 improves from 0.8$\rightarrow$27.6during learning with Naive+, a significant, but modest improvement. We ran additional experiments to study the presence of learning with abstract labels, but without reward signals (i.e., to measure the domain effect). We see no improvement over zero-shot performance, confirming the reward signal is what drives learning. [^8] Because these experiments showed no learning effect at all, we are omitting them from our figures.
Bigger Models Are Better at ICRL
We evaluate Qwen Instruct models ranging from 500M to 72B parameters using both Naive+ and Stochastic to characterize the scaling trends of ICRL. 5(a) shows the results. For all model sizes, performance improves substantially over zero-shot accuracy (measured at the first time step). However, smaller models tend to plateau at lower accuracies compared to larger models, indicating that ICRL benefits from model scale, similar to other LLM behaviors.
Stochastic is More Stable Than Naive+
An important differentiating factor between Stochastic and Naive+ is stability. The results so far (Figures 2, 4, and 5(a)) often show sudden, even if temporary dips in performance with Naive+. This instability is undesirable, because it means the performance of the model in interactions (i.e., with users) temporarily deteriorates significantly. In contrast, visually, the learning trends of Stochastic are more stable.
We quantify stability as Spearman’s rank correlation ($\rho$) between accuracy and time step. [^9] A perfect positive correlation ($\rho=1$) indicates strictly increasing accuracy over time, whereas $\rho=-1$ means performance strictly decreases. 5(b) shows the relation between stability and model sizes for all Qwen models for the three methods. Except for TREC and TREC-fine, which give inconclusive results, Stochastic exhibits more stable learning than Naive+ on Banking77, CLINC150, and NLU across all model scales. As expected, in general, larger models show higher stability, mirroring the trends in 5(a). We hypothesize that Stochastic is less sensitive to short-term fluctuations, as it relies on a smaller but more diverse set of episodes at each step.
5 Related Work
Supervised ICL
ICL was first demonstrated by Brown et al. (2020), and since then its causes (Chan et al., 2022; Xie et al., 2022; Olsson et al., 2022; Garg et al., 2022; Von Oswald et al., 2023; Hendel et al., 2023; Wang et al., 2023) and the level of learning it displays (Min et al., 2022; Lyu et al., 2023) have been studied extensively. By now, it is well established that LLMs can learn new tasks in context (Garg et al., 2022; Wei et al., 2023; Pan et al., 2023; Kossen et al., 2024; Li et al., 2024). Our work builds on this line of work, and provides the first evidence that LLMs have the innate capability to perform RL in context, and not only supervised learning (i.e., the standard way it is done), in the contextual bandit setting.
Our study would not be possible without recent increases in the context window length of LLMs (Llama Team, 2024; Abdin et al., 2024; Gemini Team, 2024). Recent work showed that model performance can continue to increase when including hundreds or thousands of ICL demonstrations (Bertsch et al., 2024; Agarwal et al., 2024). We find similar results: LLMs can continually improve when learning through ICRL until their context does not saturate. Interestingly, while some work (Zhang et al., 2024b; Mo et al., 2024; Shinn et al., 2023) find that models can learn from mistakes, we do not observe effective learning from episodes with negative rewards. It is possible that models can learn from mistakes only when explicitly reasoning (Kojima et al., 2022; Wei et al., 2022) about them (Shinn et al., 2023; Zhang et al., 2024b), but not implicitly. This is an important direction for further study.
ICRL
Likely the closest work to ours is Krishnamurthy et al.’s (2024) study of whether LLMs can solve multi-armed bandit problems, a state-less simpler RL setting than the one we are focused on. We observe similar issues to their findings with the Naive approach. They present a set of negative results, and finally are able to elicit effective learning, but through a prompting strategy that cannot generalize beyond their very simple scenario. We address this challenge by showing the strong performances of both Naive+, which includes only positive outcomes and an increased sampling temperature, and Stochastic, which features stochasticity in the prompt construction. Concurrent to our work, Nie et al. (2024) studied the contextual bandit ICRL, similar to our study. They also propose a working solution, but take a very different approach by externally tracking learning statistics commonly used in the UCB bandit learning algorithm (Auer et al., 2002), and fine-tuning or prompting a model to leverage them. In contrast, the approaches we discuss do not rely on explicitly tracking statistics or fine-tuning. Instead, we are intered in innate abilities.
Wu et al. (2024) propose benchmarks that include a simplified multi-armed bandit problem. Their baseline results with a method similar to Naive show mixed results in a setting that is even simpler than that of Krishnamurthy et al. (2024).
A few studies have also considered multi-step RL toy settings (Brooks et al., 2023; Mirchandani et al., 2023). Mirchandani et al. (2023) prompt models to improve past trajectories, and Brooks et al. (2023) simulate policy iteration with LLMs. Both works find that models cannot learn in general. Interestingly, Mirchandani et al. (2023) attribute this failure to LLMs inability to explore and find optimal solutions, as we also observed in our analysis of Naive.
Transformers and RL
Another related line of research is that of Transformers trained to solve sequential decision-making problems (Janner et al., 2021; Chen et al., 2021; Xu et al., 2022; Laskin et al., 2022; Zheng et al., 2022; Lee et al., 2023; Grigsby et al., 2024; Raparthy et al., 2024). In all these cases, Transformers (Vaswani et al., 2017) are trained from scratch. Our focus is different: we study ICRL that emerges from the process of training LLMs, without fine-tuning the LLM for this purpose.
6 Discussion and Limitations
We study the innate capabilities of off-the-shelf LLMs to perform ICRL in the contextual bandit setting. We outline a straightforward algorithm to show this behavior, and propose an enhanced version featuring stochasticity in the prompt construction, while increasing stability. We characterize ICRL, including scaling effects, stability, the importance of the reward signal, and the impact of abstract labels (i.e., that contain no semantic information).
While our work provides a plethora of insights into ICRL behavior, much remain to be studied. We intentionally selected the contextual bandit setting and classification benchmarks and focused on binary rewards to simplify the experiments and evaluation in this early stage of studying ICRL. This is a limitation of this study, and leaves open the question of applicability to more complex RL problems, where rewards are more nuanced, or where interactions comprise multiple steps. For example, math and coding tasks often require multiple steps, but also introduce complex evaluation challenges. We believe our study enables future work to study these challenges, and that this is an important direction.
The limited capability we identified to implicitly reason about episodes with negative rewards in current LLMs also outlines an important research question for future work. Our work also leaves open questions about the use of computational resources. ICRL is relatively compute-intensive, especially after the learner observes many episodes. We propose Approximate ICRL in Appendix B.3 to reduce certain forms of computational overhead, and show how it allows to trade-off compute for robustness. Further reducing computational demands is an important direction for future work.
We hope our work helps to shed light on the capabilities of contemporary LLMs, and that it lays out the ground for extensive future work, both in research and practice.
Acknowledgments
We thank Yair Feldman for proposing Spearman’s rank correlation as a stability metric, and Mustafa Omer Gul, Yair Feldman, Yilun Hua, and Robert West for insightful discussion and feedback.
This research was supported by NSF under grants No. 1750499 and OAC-2311521, NASA under award No. 20-OSTFL20-0053, a gift from Open Philanthropy, a gift from Apple, the National Artificial Intelligence Research Resource (NAIRR) Pilot, the Frontera supercomputer supported by the National Science Foundation (award NSF-OAC 1818253) at the Texas Advanced Computing Center (TACC) at The University of Texas at Austin, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award NSF-OAC 2005572). We thank Google for enabling experiments with Gemini through a gift. AB gratefully acknowledges the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, and the Allen Institute for AI. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation, NASA, or the other funders.
Appendix A Evaluation Measures in the Appendix
In the appendix, in multiple cases, we also report regret, the forgone utility from an actual model prediction in comparison to the oracle choice. Intuitively, regret measures how many interactions the model handled poorly throughout the experiment. In our experiments, regret is the accumulated number of incorrect examples throughout learning. Regret gives a single number that considers both the final performance and how fast the model reached it. A good system would reach high performance as fast as possible, making fewer mistakes overall (i.e., would have a low regret).
In some cases, we also report train accuracy as the running mean accuracy over the most recent 256 episodes.
Appendix B Additional Method Analysis
B.1 Naive and Naive+ ICRL Hyperparameters
We study the effect of the model sampling temperature $T$ on both Naive and Naive+ ICRL (Figure 6). For Naive (6(a)), we observe that varying $T$ does not significantly affect performance, and all values lead to relatively poor results. In contrast, Naive+ ICRL is highly sensitive to $T$ (6(b)): while higher temperatures can sometimes reach stronger performance, they also introduce substantial instability. Low temperatures are more stable but plateau at lower levels of accuracy. Overall, we find that $T=2.0$ achieves both good performance and stability. We adopt this value for all subsequent Naive+ experiments, including the ablations reported in Figure 3.
A related concern involves zero-shot performance (i.e., performance at time step 0). Because we use $T=1.0$ for Stochastic (and Naive) and $T=2.0$ for Naive+, it is unclear whether to measure zero-shot performance with $T=1.0$ or $T=2.0$. To ensure fairness, in all experiments combining Naive+ and Stochastic, we report the higher of these two zero-shot accuracies as our baseline. In particular, in many instances, because of this choice, the difference between final and initial performance exceeds the difference between final and zero-shot performance.
(a) Temperature Sensitivity Analysis for Naive.
(b) Temperature Sensitivity Analysis for Naive+.
Figure 6: Temperature Sensitivity Analysis for Naive and Naive+. We plot the performance of each approach across different sampling temperatures $T$. (a) For Naive, varying $T$ has little impact, and all temperature settings result in relatively poor performance. (b) Naive+ exhibits significant variability: higher $T$ values can lead to strong performance but are less stable, whereas lower $T$ values yield more stable results but with lower peak accuracy. We choose $T=2.0$ for Naive+ to achieve both strong performance and stability.
B.2 Stochastic ICRL
B.2.1 Downsampling Strategies
In our formulation of Stochastic ICRL, we downsample too large contexts by randomly removing selected episodes until they fit the model context. However, we design three strategies in total to downsample the context if we reach the limit of the LLM context window: (a) unbiased (the default strategy): randomly remove episodes from $C^{(t)}$ until it fits the context window; (b) start-biased: use the longest possible prefix of episodes from $C^{(t)}$ such that it fits the LLM context size; and (c) end-biased: use the longest possible suffix. Unbiased corresponds to the approach used in the main paper.
In practice, we never saturate the LLM context window when using Stochastic ICRL with $p_{\text{keep}}=0.1$ because our context windows are more than 100k. We conduct experiments to evaluate the above strategiesby limiting the context window of Llama to 4k or 8k tokens. Generally, we observe that start-biased strategy outperforms unbiased, which in turn performs better than end-biased, in all cases, although by only small margins. Given these results, we focus on unbiased as the most straightforward approach. Figure 8 shows the results of this analysis for Banking77 and CLINC150.
B.2.2 Hyperparameter Tuning and Sensitivity
Stochasticity in context generation is one of the key components that contribute to both Stochastic performance. It is controlled by setting $p_{\text{keep}}$. Figure 7 shows the sensitivity of Stochastic to different values of $p_{\text{keep}}$. Without stochasticity ($p_{\text{keep}}=1.0$), ICRL struggles on both models—particularly on Phi —while setting $p_{\text{keep}}$ too low retains too few examples in the context and hurts performance. Setting $p_{\text{keep}}=0.1$ strikes a good balance, yielding strong results while keeping the context short (and therefore faster to run). We fix $p_{\text{keep}}$ to 0.1 for all subsequent Stochastic experiments.
We do not optimize the sampling temperature $T$ for Stochastic in this work and simply fix it to a standard value of 1.0. It is possible that performance could improve further with a more optimal temperature selection. We leave this investigation to future work.
Figure 7: Sensitivity to $p_{\text{keep}}$ in Stochastic ICRL. We compare performance with different values of $p_{\text{keep}}$. Intermediate values learn better for both Llama and Phi.
Figure 8: Varying Maximum Context in Llama for Banking77 and CLINC150. Comparison of test accuracy and regret of Llama under varying context lengths and subsampling strategies for both Banking77 and CLINC150 datasets. Longer contexts generally enhance performance, with subtle differences observed between subsampling strategies. The difference between the strategies is negligible.
Algorithm 3 Approximate ICRL
0:
0:  Everything from Algorithm 2
0:  $K$: Number of contexts to maintain
0:
1:  Init empty contexts $\mathcal{C}\leftarrow\{[\,]^{(1)},\dots,[\,]^{(K)}\}$
2:  for $t=1,2,3,\dots$ do
3:    Sample context uniformally $C\sim\mathcal{U}\left(\mathcal{C}\right)$
4:    Observe input $x^{(t)}\sim\mathcal{D}$
5:    Sample prediction $\hat{y}^{(t)}\sim\pi(\cdot|C,x^{(t)})$
6:    Observe reward $r^{(t)}\sim R(x^{(t)},\hat{y}^{(t)})$
7:    if $r>0$ then
8:      for $k=1$ to $K$ do
9:         $b\sim\text{Bernoulli}(p_{\text{keep}})$
10:         if $b=1$ then
11:           Add episode to cached context              $\mathcal{C}[k]\mathrel{+}=(x^{(t)},\hat{y}^{(t)},r^{(t)})$
12:         end if
13:      end for
14:    end if
15:  end for
B.3 Approximate ICRL
B.3.1 Stochastic ICRL Computational Costs
An important technical difference between the Naive and Naive+ approaches and Stochastic is that, until the context window is not saturated, Naive approaches can potentially re-use past computations from caching. This is not possible in Stochastic, because each episode requires the construction of a fresh context $C^{(t)}$. The probability of encountering the same context twice, or even the same prefix, is exceptionally low even after a few episodes. This means that the context has to be computed from scratch for each input. [^10] In low-memory setups, this does not lead to noticeable slowdowns (as efficient caching would not be possible), and Stochastic can be much faster given that each context contains only $p_{\text{keep}}$of the episodes that Naive+ would use. In our setup, we empirically find this to be the case and Stochastic is significantly faster in practice.
B.3.2 Method
We propose an approximation of Stochastic ICRL that balances between computational cost and learning effectiveness. Similar to both Naive+ and Stochastic, the approximate version also excludes episodes with negative reward and, like Stochastic, focuses on exploration by stochasticity in the context.
Algorithm 3 describes Approximate ICRL. The core idea behind the approximation is to persistently store a limited number of contexts, so we can simply gradually expand them with new episodes, rather than always create and compute new contexts. We maintain $K$ contexts $\mathcal{C}$, which all start empty (line 1). At each time step $t$, we sample a context $C$ from the $K$ contexts (line 3), and use it for episode $t$ (lines 46. If the reward $r^{(t)}>0$, we use the episode to expand all contexts stochastically. For each context in $\mathcal{C}$, we expand it with the $t$-th episode with a probability of $p_{\text{keep}}$ (lines 811).
Approximate introduces stochasticity in two places: sampling the context to use for each episode and the expansion of the stored contexts. In Algorithm 3, we use uniform sampling to choose the context (line 3). This is a uniform approximation of the probability of a context, which can also be easily computed exactly using the probabilities of the episodes it contains and $p_{\text{keep}}$. In practice, we find the exact computation to work poorly, because contexts that are assigned more episodes or have low probability episodes quickly receive very low probability, and are not used. 10(b) shows this experimental analysis. We use uniform sampling throughout our experiments.
The level of approximation the algorithm provides depends on the resources available. For example, one can allocate each context to a compute unit, so a machine with eight compute units (e.g., GPUs) will support $K=8$. Approximate is a strict approximation of Stochastic in the sense that coupling the exact context sampling strategy with $K\to\infty$ gives Stochastic. However, the approximation is limited in handling contexts that extend beyond the LLM window length. Overcoming this while maintaining the efficiency of the approximation is an important direction for future work.
B.3.3 Results
We test Approximate on Llama and Phi only, and show the results in Figure 9. If not specified, we choose $K=8$ for Approximate.
Figure 9: Performance of ICRL. Stochastic and Approximate held-out test results for both Llama and Phi and all semantic-labels tasks.
Approximate is an Effective Alternative to Stochastic
In Figure 9, Approximate performs almost as well as Stochastic ICRL when using Llama, across all tasks. The results are very different with Phi: despite early learning, Approximate deteriorates quickly. This stems from one of the contexts being biased towards one label and therefore predicting only this label. Eventually, the bias towards the label spreads to other contexts, leading to the collapse in performance we observe. It is empirically possible to recover, as we see in Banking77 later in the experiment, but the chance of it happening seems low. The success of Llama and failure of Phi with $K=8$ show that different LLMs have different sensitivity to the approximation. 10(a) shows that that with a higher number of contexts $K>32$ Phi is able to effectively learn, indicating Phi needs a higher computational budget. On the other hand, Llama is robust to the approximation, with most values performing similarly to Stochastic, except with the lowest values of $K$.
Approximate Reduces Compute Needs.
We measure the reduction of tokens processed in Approximate compared to Stochastic throughout full ICRL runs. We approximate this measure by computing at each step the number of tokens required for a forward call and subtracting the number of tokens of the sequence with the longest common prefix processed in a previous step, as it would be possible to use the KV cache for all the tokens in the common prefix (assuming infinite memory). We find that Stochastic processes two orders of magnitude more tokens than Approximate. Table 7 provides numerical results for this analysis.
(a) Effect of the number of contexts.
(b) Effect of the context sampling strategies.
Figure 10: Comparison of Approximate parameters. (a) Effect of the number of contexts $K$. We report test accuracy for Llama and Phi. Phi proves more sensitive to this approximation. Generation degenerates for low $K$, while the model can learn for $K\geq 32$. Llama can learn with all $K$, although higher values perform better. (b) Comparison of exact and uniform sampling. We report test accuracy at the final step and regret for Llama. Uniform sampling strategy is consistently better.
Appendix C Experimental Setup
We conduct experiments on various type of GPUs: 40GB A100, 80GB A100, 80GB H100, 48GB A6000. For experiments with 70B and 32B models, we use 4 80GB A100/H100 or 8 48GB A6000. For experiments with 14B models, we use 2 80GB A100/H100. For experiments with 7B or smaller models, we use 1 80GB A100/H100 or 2 48GB A6000 / 40GB A100. For efficient inference, we use vllm (Kwon et al., 2023).
C.1 Prompt Design
We report prompt examples from ICL (Figure 11) and ICRL (Figure 12) experiments. We show the prompts for Llama as an example. In all cases, we show the prompts with two in-context examples.
[…]
Figure 11: An example of prompt of ICL for Llama.
[…]
Figure 12: An example of prompt of ICRL for Llama.
C.2 Context Windows and Episode Capacity
For each task and model combination, we conservatively estimate the maximum number of examples that could fit within the context window. This is done by including all observed examples in descending order of token count in the prompt, assuming the model consistently responds with the longest label and that the formatted reward message is at its maximum length. We perform this calculation using the maximum context window for all models. Additionally, for Llama, we repeat the process with context windows of 4,096 and 8,192 tokens specifically for the Banking77 and CLINC150 tasks. Table 1 reports episode capacity.
Table 1: Maximum number of episodes supported by model and task, given a specific context window. We compute the maximum number of episodes supported by the context window of Llama, Phi, Qwen, and Gemini across all tasks, including 4k and 8k tokens for Llama, with Banking77 and CLINC150 only.
Phi Llama Qwen Gemini
Task 128k tokens 4k tokens 8k tokens 128k tokens 128k tokens 1M tokens
Banking77 1538 34 74 1673 1672 -
CLINC150 2241 60 126 2384 2184 -
NLU 2397 - - 2425 2424 -
TREC 2848 - - 2919 2896 -
TREC-fine 2584 - - 2776 2755 -
Abs. Banking77 - - - 1924 1788 13007
Abs. CLINC150 - - - 2485 2270 19501
Abs. NLU - - - 2475 2285 24603
Abs. TREC - - - 2529 2308 5953
Abs. TREC-fine - - - 2531 2308 5953
C.3 Datasets
We use 5 classification tasks (and the corresponding abstract-label variants) in our experiments:
  • Banking77 (77 labels; Casanueva et al. (2020)). It involves 77 labels and aims to detect the intent of user queries in an economic context. For example, one label could be “balance not updated after cheque or cash deposit”.
  • CLINC150 (150 labels; Larson et al. (2019). It includes 150 labels, also focusing on intent classification. An example label is “calendar update”. While the original dataset was designed to detect out-of-scope queries, we concentrate solely on classifying the 150 defined intents, excluding out-of-scope queries from our analysis.
  • NLU (68 labels; Liu et al. (2021)). This dataset includes queries grouped in 68 unique categories for human-robot interaction in home domain (for example, one label is “audio volume mute”.
  • TREC and TREC-fine (respectively 6 and 50 labels; Li & Roth (2002); Hovy et al. (2001)). Both are question classification dataset where the goal is to classify the type of question. Each example contains both a fine label (that we use in TREC-fine), as “entity vehicle”, and a coarse one (used in TREC), as “entity”. TREC-fine includes 50 categories, while TREC groups them in only 6 categories.
All of these datasets are challenging because of the big number of different labels, and the sometimes subtle differences between labels. Moreover, in our setting we do not provide any information about the list of potential labels (except for the “With Exemplars” abstract-label experiments), challenging the model to either follow previously discovered labels or try to find new, more suitable ones (i.e., exploitation vs exploration – Sutton & Barto (2018)).
Appendix D Additional Results
Table 2: Detailed Figures for Figure 2. We report three key figures for each dataset and method: initial (zero-shot) accuracy, final (post-ICRL) accuracy, and regret (total mistakes). (a) contains the results for Llama, (b) for Qwen.
Naive Naive+ Stochastic Upper Bound
Dataset 0-step Final Reg. 0-step Final Reg. 0-step Final Reg. Acc.
Acc. Acc. Acc. Acc. Acc. Acc.
Banking77 0.172 0.104 8755 0.152 0.648 4358 0.172 0.660 4613 0.846
CLINC150 0.246 0.292 7126 0.092 0.836 2280 0.246 0.814 2790 0.944
NLU 0.338 0.322 6868 0.286 0.578 4486 0.338 0.706 3545 0.844
TREC 0.476 0.412 3692 0.326 0.744 2235 0.476 0.836 1183 0.872
TREC-fine 0.238 0.29 4390 0.070 0.610 2585 0.238 0.740 2183 0.760
Naive Naive+ Stochastic Upper Bound
Dataset 0-step Final Reg. 0-step Final Reg. 0-step Final Reg. Acc.
Acc. Acc. Acc. Acc. Acc. Acc.
Banking77 0.064 0.174 8265 0.070 0.572 4433 0.062 0.722 4274 0.862
CLINC150 0.154 0.108 8691 0.138 0.802 2378 0.154 0.838 2913 0.960
NLU 0.216 0.230 7369 0.248 0.788 2967 0.208 0.748 3235 0.874
TREC 0.430 0.292 3883 0.372 0.664 2477 0.440 0.806 1181 0.880
TREC-fine 0.146 0.014 4734 0.122 0.536 3337 0.144 0.704 1920 0.812
Table 3: Detailed Metrics for Figure 3. We report three key metrics for each dataset and method: initial (zero-shot) accuracy, final (post-ICRL) accuracy, and regret (total mistakes). (a) shows results for Banking77, (b) for CLINC150.
(a) Banking77
Naive Stochastic
Reward 0-step Final Reg. 0-step Final Reg.
Acc. Acc. Acc. Acc.
None 0.156 0.034 9426 0.172 0.308 7132
Only neg. 0.152 0.060 9331 0.172 0.214 7814
Only pos. 0.152 0.648 4358 0.172 0.660 4613
Both pos. and neg. 0.156 0.184 8624 0.172 0.458 5943
Noisy pos. and neg. 0.156 0.174 8713 0.172 0.394 6256
Inv. pos. and neg. 0.152 0.098 9234 0.172 0.282 7047
(b) CLINC150
Naive Stochastic
Reward 0-step Final Reg. 0-step Final Reg.
Acc. Acc. Acc. Acc.
None 0.092 0.056 9214 0.246 0.388 6555
Only neg. 0.092 0.082 9010 0.246 0.208 7496
Only pos. 0.092 0.836 2280 0.246 0.814 2790
Both pos. and neg. 0.092 0.170 8280 0.246 0.582 4688
Noisy pos. and neg. 0.092 0.146 8454 0.246 0.586 4810
Inv. pos. and neg. 0.092 0.098 8995 0.246 0.448 5865
Table 4: Detailed Figures for Figure 4. We report three key figures for each dataset and method: initial (zero-shot) accuracy, final (post-ICRL) accuracy, and regret (total mistakes). (a) contains the results for Llama, (b) for Qwen, and (c) for Gemini.
Abs. Banking77 Abs. CLINC150 Abs. NLU Abs. TREC Abs. TREC-fine
Method 0-step Final Reg. 0-step Final Reg. 0-step Final Reg. 0-step Final Reg. 0-step Final Reg.
Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc.
Naive 0.020 0.018 9909 0.002 0.000 9928 0.010 0.028 9674 0.030 0.086 4032 0.034 0.030 4828
Naive+ (w/o Ex.) 0.020 0.454 7164 0.008 0.276 8507 0.010 0.476 6789 0.118 0.790 2046 0.080 0.144 4597
Naive+ (w/ Ex.) 0.024 0.178 4704 0.000 0.590 5727 0.056 0.656 4166 0.238 0.792 2214 0.006 0.196 4153
Stochastic (w/o Ex.) 0.018 0.350 8280 0.002 0.162 9437 0.010 0.428 7288 0.030 0.838 2244 0.034 0.258 4767
Stochastic (w/ Ex.) 0.030 0.584 4380 0.006 0.650 6276 0.162 0.722 3725 0.238 0.840 2424 0.008 0.036 4773
Up. Bound Acc. 0.780 0.886 0.750 0.886 0.612
Abs. Banking77 Abs. CLINC150 Abs. NLU Abs. TREC Abs. TREC-fine
Method 0-step Final Reg. 0-step Final Reg. 0-step Final Reg. 0-step Final Reg. 0-step Final Reg.
Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc.
Naive 0.016 0.002 9877 0.012 0.014 9892 0.010 0.014 9731 0.124 0.13 3876 0.010 0.024 4810
Naive+ (w/o Ex.) 0.014 0.442 7391 0.008 0.204 8824 0.014 0.622 6479 0.136 0.526 3304 0.010 0.146 3713
Naive+ (w/ Ex.) 0.248 0.626 4618 0.048 0.602 4481 0.162 0.656 3818 0.158 0.776 1564 0.004 0.526 3044
Stochastic (w/o Ex.) 0.016 0.208 8653 0.012 0.134 9449 0.010 0.564 6128 0.124 0.886 1591 0.010 0.532 3868
Stochastic (w/ Ex.) 0.316 0.680 4353 0.080 0.712 4385 0.182 0.706 3882 0.200 0.890 1520 0.002 0.462 4440
Up. Bound Acc. 0.844 0.902 0.822 0.892 0.574
Abs. Banking77 Abs. CLINC150 Abs. NLU Abs. TREC Abs. TREC-fine
Method 0-step Final Reg. 0-step Final Reg. 0-step Final Reg. 0-step Final Reg. 0-step Final Reg.
Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc. Acc.
Naive 0.018 0.000 9996 0.008 0.010 9929 0.048 0.004 9938 0.076 0.012 4841 0.002 0.018 4915
Naive+ (w/o Ex.) 0.014 0.174 8858 0.002 0.144 9264 0.044 0.218 7978 0.104 0.900 1308 0.006 0.130 3998
Naive+ (w/ Ex.) 0.430 0.778 2776 0.356 0.818 1711 0.474 0.774 2582 0.480 0.896 1127 0.134 0.682 2448
Stochastic (w/o Ex.) 0.014 0.214 8625 0.008 0.060 9710 0.050 0.290 8281 0.072 0.876 1505 0.002 0.060 4920
Stochastic (w/ Ex.) 0.440 0.792 3292 0.434 0.858 1984 0.480 0.794 2560 0.494 0.872 1439 0.138 0.636 3420
Up. Bound Acc. 0.902 0.964 0.890 0.946 0.860
Table 5: Detailed Metrics for 5(a). We report three key metrics for each dataset and method: initial (zero-shot) accuracy, final (post-ICRL) accuracy, and regret (total mistakes). (a)–(g) show results for different sizes of Qwen2.5.
(a) Qwen2.5 500M
Naive+ Stochastic
Dataset 0-step Final Reg. 0-step Final Reg.
Banking77 0.082 0.284 7351 0.146 0.364 6874
CLINC150 0.062 0.404 5959 0.118 0.552 5157
NLU 0.088 0.358 6746 0.146 0.544 6162
TREC 0.290 0.382 2260 0.318 0.514 2182
TREC-fine 0.032 0.154 3955 0.034 0.308 3961
(b) Qwen2.5 1.5B
Naive+ Stochastic
Dataset 0-step Final Reg. 0-step Final Reg.
Banking77 0.074 0.258 6069 0.092 0.524 5824
CLINC150 0.066 0.482 4660 0.182 0.684 3945
NLU 0.242 0.302 5084 0.288 0.636 4619
TREC 0.274 0.302 2150 0.362 0.754 1835
TREC-fine 0.042 0.538 2979 0.074 0.680 2681
(c) Qwen2.5 3B
Naive+ Stochastic
Dataset 0-step Final Reg. 0-step Final Reg.
Banking77 0.082 0.532 4959 0.094 0.614 4852
CLINC150 0.156 0.778 2370 0.148 0.812 2936
NLU 0.270 0.734 3355 0.266 0.772 2666
TREC 0.428 0.744 2482 0.506 0.730 1246
TREC-fine 0.098 0.520 3439 0.214 0.620 2131
(d) Qwen2.5 7B
Naive+ Stochastic
Dataset 0-step Final Reg. 0-step Final Reg.
Banking77 0.070 0.572 4433 0.062 0.722 4274
CLINC150 0.138 0.802 2378 0.154 0.838 2913
NLU 0.248 0.788 2967 0.208 0.748 3235
TREC 0.372 0.664 2477 0.440 0.806 1181
TREC-fine 0.122 0.536 3337 0.144 0.704 1920
(e) Qwen2.5 14B
Naive+ Stochastic
Dataset 0-step Final Reg. 0-step Final Reg.
Banking77 0.126 0.506 4902 0.144 0.730 4136
CLINC150 0.192 0.800 2505 0.248 0.882 2439
NLU 0.348 0.574 4198 0.376 0.748 3228
TREC 0.492 0.786 2105 0.556 0.904 919
TREC-fine 0.252 0.522 2989 0.322 0.734 1869
(f) Qwen2.5 32B
Naive+ Stochastic
Dataset 0-step Final Reg. 0-step Final Reg.
Banking77 0.132 0.518 5015 0.144 0.778 3733
CLINC150 0.220 0.838 2164 0.280 0.884 2467
NLU 0.360 0.702 3588 0.380 0.768 2842
TREC 0.590 0.622 2185 0.646 0.920 770
TREC-fine 0.264 0.516 2800 0.354 0.662 1599
(g) Qwen2.5 72B
Naive+ Stochastic
Dataset 0-step Final Reg. 0-step Final Reg.
Banking77 0.186 0.592 4404 0.212 0.794 3512
CLINC150 0.328 0.870 1714 0.360 0.906 1983
NLU 0.380 0.776 2520 0.438 0.794 2609
TREC 0.392 0.656 2277 0.416 0.898 782
TREC-fine 0.100 0.706 2398 0.170 0.764 1614
Table 6: Detailed Stability Metric $\rho$ for 5(b). Each cell contains the stability metric $\rho$ for the corresponding dataset and method. (a)–(g) show results for different sizes of Qwen2.5.
(a) Qwen2.5 500M
Naive Naive+ Stochastic
Banking77 -0.343 0.491 0.710
CLINC150 -0.459 0.634 0.877
NLU -0.025 0.045 0.868
TREC -0.500 0.264 0.700
TREC-fine -0.202 0.164 0.724
(b) Qwen2.5 1.5B
Naive Naive+ Stochastic
Banking77 0.209 0.427 0.875
CLINC150 -0.369 0.278 0.949
NLU -0.260 0.088 0.851
TREC -0.500 0.064 0.501
TREC-fine 0.500 0.882 0.697
(c) Qwen2.5 3B
Naive Naive+ Stochastic
Banking77 -0.069 0.710 0.932
CLINC150 -0.652 0.748 0.956
NLU -0.857 0.711 0.925
TREC -0.500 0.882 0.773
TREC-fine 0.500 0.210 0.564
(d) Qwen2.5 7B
Naive Naive+ Stochastic
Banking77 -0.653 0.694 0.989
CLINC150 -0.794 0.679 0.985
NLU -0.570 0.661 0.963
TREC -0.230 -0.134 0.314
TREC-fine -0.674 0.784 0.487
(e) Qwen2.5 14B
Naive Naive+ Stochastic
Banking77 0.394 0.548 0.964
CLINC150 -0.730 0.956 0.981
NLU -0.494 0.337 0.859
TREC -0.790 0.927 0.441
TREC-fine -0.691 0.540 0.825
(f) Qwen2.5 32B
Naive Naive+ Stochastic
Banking77 -0.218 0.684 0.972
CLINC150 -0.419 0.939 0.989
NLU 0.035 0.887 0.945
TREC -0.791 0.074 0.822
TREC-fine -0.849 0.886 0.292
(g) Qwen2.5 72B
Naive Naive+ Stochastic
Banking77 -0.456 0.894 0.965
CLINC150 0.159 0.854 0.986
NLU -0.633 0.461 0.910
TREC -0.500 0.936 0.765
TREC-fine -0.500 0.970 0.600
Table 7: Tokens processed in Approximate compared to Stochastic throughout full ICRL runs. Stochastic processes two orders of magnitude more tokens than Approximate.
Phi Llama
Task Expl. Approx. Ratio Expl. Approx. Ratio
Banking77 87,369,607 510,786 171 102,282,989 539,367 190
CLINC150 105,545,002 398,677 265 122,455,599 440,019 278
NLU 89,894,548 409,680 219 114,517,653 433,254 264
TREC 29,306,971 212,855 138 34,509,170 229,046 151
TREC-fine 20,658,980 222,955 93 25,522,358 234,884 109
Figure 13: Detailed visualization of Approximate for Llama, Banking77 with exact context sampling. We report test accuracy (top left), a 256-step running average of the training accuracy (bottom left), the training accuracy of each context (top right), and the hit rate of each context (bottom right).
Figure 14: Detailed visualization of Approximate for Llama, Banking77 with uniform context sampling. We report test accuracy (top left), a 256-step running average of the training accuracy (bottom left), the training accuracy of each context (top right), and the hit rate of each context (bottom right).
Figure 15: Detailed visualization of Approximate for Llama, CLINC150 with exact context sampling. We report test accuracy (top left), a 256-step running average of the training accuracy (bottom left), the training accuracy of each context (top right), and the hit rate of each context (bottom right).
Figure 16: Detailed visualization of Approximate for Llama, CLINC150 with uniform context sampling. We report test accuracy (top left), a 256-step running average of the training accuracy (bottom left), the training accuracy of each context (top right), and the hit rate of each context (bottom right).
Figure 17: Detailed visualization of Approximate for Phi, Banking77 with uniform context sampling. We report test accuracy (top left), a 256-step running average of the training accuracy (bottom left), the training accuracy of each context (top right), and the hit rate of each context (bottom right).
Figure 18: Detailed visualization of Approximate for Phi, CLINC150 with uniform context sampling. We report test accuracy (top left), a 256-step running average of the training accuracy (bottom left), the training accuracy of each context (top right), and the hit rate of each context (bottom right).